OpenStudy (anonymous):

Solve the system by elimination: -2x + 2y + 3z = 0 -2x - y + z =-3 2x + 3y + 3z = 5

OpenStudy (shamil98):

Multiply the third equation by 2 and then add the equations together. Then you'll have a two variable system, from there eliminate z or y and solve for the corresponding variable.

OpenStudy (anonymous):

@shamil98 do you know how to find the length of the minor axis of an ellipse

OpenStudy (anonymous):

Thanks I'll try it an report back to you if i have any problems later :)

OpenStudy (anonymous):

I'm so confused I work it out an got z = 3/2 -3y/2

OpenStudy (anonymous):

Is this right or wrong?

Directrix (directrix):

(1) -2x + 2y + 3z = 0 (2) -2x - y + z =-3 (3) 2x + 3y + 3z = 5 -------------------------- Take equations (1) and (3) and add. This will eliminate the x terms in those two equations (1) -2x + 2y + 3z = 0 (3) 2x + 3y + 3z = 5 ----------------------- * 5y + 6z = 5

Directrix (directrix):

Now, take equations (2) and (3) and add them. That will eliminate the x variable. (2) -2x - y + z =-3 (3) 2x + 3y + 3z = 5 --------------------- ** 2y + 4z = 2

Directrix (directrix):

Now, take the two equations that resulted from eliminating variable x in (1) and (3) and then in (2) and (3). In those two, choose eliminate variabel y OR z. * 5y + 6z = 5 ** 2y + 4z = 2

Directrix (directrix):

* 5y + 6z = 5 ** 2y + 4z = 2 -------------- Let's eliminate variable y. To do this, multiply each member of the first equation * by 2 and the second equation ** by -5.

OpenStudy (anonymous):

z = 5/6

Directrix (directrix):

* 5y + 6z = 5 ---by 2--> 10y + 12z = 10 ** 2y + 4z = 2 --> by -5 > -10y - 20z = -10 ----------------- Add ----> -8z = 0 z = 0 @Urebaa When I multiplied equation ** by -5, I made two errors. I have corrected them here. And, I went ahead and solved for z.

Directrix (directrix):

So, z = 0.

Directrix (directrix):

Go to equation * 5y + 6z = 5 and substitute 0 in for variable z. Then, solve for y. Post what you get, okay? 5y + 6*0 = 5

OpenStudy (anonymous):

5y + 0 = 5 5y = 5 y = 1

Directrix (directrix):

Yes. Now, we need the value of x.

Directrix (directrix):

Return to any one of the three original equations: (1) -2x + 2y + 3z = 0 Substitute 0 for variable z, 1 for variable y, and solve for variable x. (1) -2x + 2*1 + 3*0 = 0 @Urebaa Solve for x and post what you get, okay?

OpenStudy (anonymous):

x = 1

Directrix (directrix):

Yes. The solution is (1, 1, 0). @Urebaa